OpenStudy (anonymous):

(12a + 4)(3a - 1) = 0

OpenStudy (anonymous):

FOIL throw the aluminum foil onto it that's how

OpenStudy (anonymous):

i don't even know what foil means let me look it up

OpenStudy (anonymous):

FOIL first, outside, inside, last FOIL

OpenStudy (anonymous):

maybe we don't even need to FOIL. really what do you want? what are you looking for? what is this i don't even know? solve for 'a' or what? aluminum foil on baked potato?

OpenStudy (dwelsher):

distribute it out 36a^2--4 (6a-2)(6A+2)=0 a=1/3,a=-1/3

OpenStudy (anonymous):

You have the product of two binomials, which equals zero: \[(12a + 4)(3a - 1) = 0\] Now for the sake of simplicity in explanation, lets call \((12a + 4) = A\), and \((3a - 1) = B\), so we have \(A * B = 0\) The product of two things is zero only if one or both of the thing are zero, so \(A = 0\) is a solution, and \(B = 0\) which gives \(12a + 4 = 0\), \(a = -{1 \over 3}\) or \(3a - 1= 0\), \(a = {1 \over 3}\)

OpenStudy (anonymous):